This problem contains three parts. In Part A, you are asked to determine the parameters of the profit function - the contribution margin and the fixed costs. You should not move on to Parts B and C until you get Part A correct, or until you use up all of your tries. The correct profit function parameters will be given to you after you have completed Part A - you should use these parameters in Parts B and C.

The problem is worth 20 points. Parts A and B are worth eight points each; Part C is worth four points. To get the eight points for Part A, you must answer both questions correctly; to get the eight points for Part B, you must answer all four questions correctly. You get eight tries on Part A, eight on Part B, and four on Part C.

______________________________________________________

ableVision has been approached by the City of Mirada to run its cable operations in 2020. After negotiating with key parties, CableVision has made the following agreements:

  • It will offer Mirada residents a basic set of 25 cable television stations at a rate of $29.99 per month.
  • It will pay the city $1,300,000 per year plus $3.75 of the monthly cable subscriber rate per month to maintain the physical facilities.
  • It will actually pay another company a monthly fixed fee of $55,000 plus $9.00 per cable subscriber per month to broadcast the 25 channels.

CableVision estimates that operating costs for billing, program news mailings, etc. will be $130,000 per month plus 8% of the monthly subscriber rate.

CableVision has several questions about its monthly revenues, costs, and profits in 2020.


REQUIRED [ROUND YOUR ANSWER TO PART A, QUESTION 1 TO THE NEAREST CENT; ROUND ALL OTHER ANSWERS TO THE NEAREST UNIT OR NEAREST DOLLAR.]

Part A (8 tries; 8 points)
1. What is the estimated monthy contribution margin per cable subscriber for CableVision in 2020?   

2. What are the estimated total monthly fixed costs for CableVision in 2020?   

Incorrect. Tries 3/8 Previous Tries


Part B (8 tries; 8 points)
1. What is CableVision's estimated monthly operating income in 2020 if 16,000 residents subscribe?   

2. How many monthly subscribers would be required for CableVision to break even in 2020?   

3. How many monthly subscribers would be required for CableVision to earn $21,000 per month in 2020?   

4. Assuming a tax rate of 30%, how many monthly subscribers would be required for CableVision to earn $21,000 per month in 2020?   

Tries 0/8


Part C (4 tries; 4 points)
Some of CableVision's managers are uncertain about their estimate of monthly fixed operating costs. Assuming that 19,000 residents subscribe, how large can monthly fixed operating costs be for CableVision to still earn $21,000 per month in 2020 (ignore taxes)?
Answer #1
Part A
1) Estimated Monthly Contribution Margin per cable subscriber
Particulars Amount Amount
Sales Revenue(Subcriber's Rate) $29.99
Less : Variable Cost per unit
Maintenance Cost $3.75
Broadcasting Fee $9.00
Operating Cost ($ 29.99*8%) $2.40
$15.15
Contribution per unit $14.84
2) Monthly Fixed Cost
Particulars Amount
Charges for maintenance of Physical Facility ($ 1,300,000/12) $108,333
Monthly Broadcast Fee $ 55,000
Operating Costs $130,000
Monthly Fixed Cost $293,333
Part B
1) Monthly Operating Income
Number of Subscribers 16000
Contribution per subscriber $14.84
Total Contribution (Subs*Cont.) $237,453
Less: Fixed cost $293,333
Monthly Operating Loss -$ 55,881
2) Monthly Subscribers to Break Even
BEP ( in units) = Fixed cost /Contribution per Subs.
BEP = $ 293,333 / $14.84
BEP = 19767 Subscribers (Rounded off)
3) Monthly Subscribers to earn $ 21,000
Number of Subs. Required = (Fixed Cost + Profit)/Contribution per Subs.
Number of Subs. Required = (293333 + 21000)/ $14.84
Number of Subs. Required = 21182 Subscribers (Rounded off)
4) Monthly Subscribers to earn $ 21,000(Tax@30%)
Profit required (After Tax) = $ 21,000
Profit Pre Tax = 21,000 *100/70 (100-30)
Profit Pre Tax = $ 30,000
Number of Subs. Required = (Fixed Cost + Profit)/Contribution per Subs.
Number of Subs. Required = (293333 + 30000)/ $14.84
Number of Subs. Required = 21788 Subscribers (Rounded off)
Verification :
Particulars Amount
Sales Revenue ( 21788 *$29.99) $653,422
Variable cost ( 21788 * $15.15) $330,089
Contribution $323,333
Les: Fixed Cost $293,333
Operating income $ 30,000
Tax @ 30% $ 9,000
Net income $ 21,000
Part C Maximum Monthly Fixed Cost
Contribution Per Subscriber $14.84
Profit Required $21,000
Fixed cost $163333 + X
Number of Subscribers 19000
Total Contribution (19000*$14.84) $281,960
Now,
Total Contribution - Fixed Cost = $21,000
$ 281,960 - ($163,333+ x) = $ 21,000
281960 - 163333 - x = 21,000
118627 -x = 21000
Operating Fixed Cost = $ 97627
Therefore, the Operating Fixed Cost should not exceed $ 97,627 to still earn an income of $ 21,000 per month.